OpenStudy (anonymous):

If a wheel that has a radius of 1 foot rolls five complete revolutions, how far will it travel? Use pi = 3.14

OpenStudy (anonymous):

This is the picture

OpenStudy (anonymous):

@jdoe0001 @Abmon98 CAN YOU HELP ME WHEN YOU GET THE TIME AND CHANCE??

OpenStudy (abmon98):

i think we should use the circumference of a circle and multiply by 5 as it says 5 revolutions

OpenStudy (anonymous):

What do you think @Glorenda49

OpenStudy (abmon98):

(2*pi*1)*5=10pi

OpenStudy (anonymous):

The answer is 31.4

OpenStudy (anonymous):

i took what you told me when you said multiply the circumference of the circle by 5 @Abmon98

OpenStudy (anonymous):

he's answered the question for you

OpenStudy (anonymous):

He didn't answer it he just showed how 2 get it

OpenStudy (anonymous):

Circumference of a circle C = 2 π r C = 2 * 3.14 * 1 C= 6.28 I believe that you'll multiply the circumference 6.28 by 5 and you'll get 31.4

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (anonymous):

You're welcome :)